OpenStudy (waheguru):

Marcel has $40 to spend on amusement park rides. Tickets cost $1.50 without a special membership pass, or $1.25 with a membership pass. A membership pass costs $5.00. Should Marcel buy a membership pass? Use mathematical reasoning to justify your answer. My answer was he should buy membership which will allow him to get 28 rides Can you show me your equations to solve this, I wanna see if they match. Thanks

OpenStudy (anonymous):

Alright, so, the number of rides without a membership pass will be given by this, where y is the number of rides and x is the money he has: y = x/1.5 With a membership pass : y = (x-5)/1.25 If you substitute x with 40 in both cases, you can get the number of rides

OpenStudy (waheguru):

Am I right?

OpenStudy (anonymous):

yep, with membership, you have 28 rides and without, you have 26 and some change left

OpenStudy (waheguru):

Thanks -